OpenStudy (solomonzelman):

I am going to do a concavity problem and abs max and min. I will just need a verification and corrections if anything. (please, patience....)

OpenStudy (surry99):

ok

OpenStudy (solomonzelman):

Given that: \(\large\color{black}{ f(x)=x^4-4x^3 }\) So, \(\large\color{black}{ f'(x)=4x^3-12x^2=4x^2(x-3) }\) \(\large\color{black}{ f''(x)=12x^2-24x=12x(x-2) }\) Critical numbers, \(\large\color{black}{ 0=4x^2(x-3) }\) \(\large\color{black}{ x=3,0 }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ f''(3)=0 }\) and \(\large\color{black}{ f'(3)>0 }\) So, \(\large\color{black}{ f(3) }\) is the local minimum.

OpenStudy (solomonzelman):

f''(x)=0, when x=0 or 2. {-inf,0) concave up (0,2) concave down (2,0) concave up

OpenStudy (solomonzelman):

the last interval should be (2,inf) and it is concave up

OpenStudy (solomonzelman):

(0,0) is an inflection point, since the concavity changes there. (From up to down)

OpenStudy (solomonzelman):

This is just a discussion question.

OpenStudy (solomonzelman):

2,-16 is another inflection point, and the concavity changes from down to up.

OpenStudy (solomonzelman):

ohh, ithe f(x) is on [0,4] My bad sorry

OpenStudy (solomonzelman):

I never post questions correctly-:(

OpenStudy (anonymous):

Take your time; just let us know when you're ready for feedback. :-)

OpenStudy (solomonzelman):

and 4 is also a critical number --- closed interval . ready.

OpenStudy (solomonzelman):

i AM DONE

OpenStudy (xapproachesinfinity):

the domain here does really matter?

OpenStudy (anonymous):

are you looking for max, min, concavity what?

OpenStudy (xapproachesinfinity):

doesn't i mean

OpenStudy (solomonzelman):

don't think the interval matters.

OpenStudy (anonymous):

i do

OpenStudy (solomonzelman):

I mean the [0,4] doesn't matter.

OpenStudy (anonymous):

ii think it does

OpenStudy (anonymous):

if you are looking for a max, you will not find one on \((-\infty,\infty)\) but you will certainly find one on \([0,4]\)

OpenStudy (solomonzelman):

for the absolute mAXIMUM it does.

OpenStudy (solomonzelman):

it would be when x=4

OpenStudy (anonymous):

you are right, if you mean not for the minimum

OpenStudy (solomonzelman):

the minimum is at x=3

OpenStudy (solomonzelman):

But have I left anything out?

OpenStudy (anonymous):

\[x^3(x-4)\] has a max of \(0\) at \(0\) and \(4\) on the closed interval

OpenStudy (solomonzelman):

max of 0?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[f(0)=f(4)=0\]

OpenStudy (solomonzelman):

Ohh, the maximum is 0 and 4

OpenStudy (solomonzelman):

I see

OpenStudy (anonymous):

oh no!

OpenStudy (anonymous):

the maximum is the output, not the input

OpenStudy (anonymous):

A couple things: I think you meant to say that \(f^{\prime}(3) = 0\) and that \(f^{\prime\prime}(3) > 0\). Your conclusion that it's a minimum is correct (via the second derivative test). We have no other critical points in the interval [0,4] since we don't allow critical points to be the endpoints of an interval. Now to determine absolute max and min, evaluate f(x) at the endpoints of the interval and at the sole critical point x=3: \[\begin{array}{c|c} x & f(x) \\\hline 0 & 0 \\ 3 & -27\\ 4 & 0 \end{array}\] Hence, the absolute max is 0 at x=0,4 and the absolute min is -27 at x=3.

OpenStudy (solomonzelman):

I mean the max is at x=0 and at x=4

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

yes, Christopher:)

OpenStudy (solomonzelman):

tnx everyone:) I have studied this second deriv. test and concavity the last lecture, but hope I'll be able to decently do these at some point.

OpenStudy (solomonzelman):

Question closed.

OpenStudy (xapproachesinfinity):

there is a nice table that gathers all the info from f' and f" critical point, points of inflaction... much better to read! like this http://commons.wikimedia.org/wiki/File:Tableau_variation_polynome_1_3_1.png

OpenStudy (solomonzelman):

My teacher prefers/requires (even if just prefers, I don't want to have a bad impression on him) it like a list of information Critical numbers: a, b, c, whatever they are Concave up and down, in interval notations in order of the domain, if you know what I mean. Like, [2,3) concave up (3,6) concave down (6,7) concave up like this in order of the x values. see what I mean? This si how I am supposed to enter my information on tests, h/w and discussions.

OpenStudy (solomonzelman):

but ty though.

OpenStudy (xapproachesinfinity):

i see... well it depends on a person after all!

OpenStudy (anonymous):

thats cool